Hofstadler Inc.’s common stock currently trades at $105.25 per share. It is expected to pay an annual dividend of $4.50 at the end of the year, and the constant growth rate is 4.0% a year.  

a. What is the company’s cost of retained earnings (internal equity)? (5 points)
b. What is the company’s cost of new stock, if flotation costs are 5%? (5 points)
Answer #1
Solution:
Cost of retained earnings (in %) 8.28 %
Working Notes:
Cost of retained earnings means cost of shareholder fund which is already in the firm, so flotation cost is not required for internal cost , it is not part of issuance cost new equity shares.
Cost of Equity (Ke) = The return shareholders are expecting=cost of retained earnings=??
Using Gordon growth model : P0 = D1 / (Ke - g),
ke = Cost of retained earnings=??
Po=current share price = $105.25 per share
g= growth rate= 4.0%
D1= Expected Dividend at t=1= $4.50 per share
P0 = D1/(Ke -g)
Ke = (D1/P0) + g
Ke = (4.50/105.25) + 4.0%
Ke = 0.0427553 + 4.0%
Ke=4.27553% + 4.0%
Ke = 8.27553%
Cost of retained earnings 8.28%
b. Cost of new stock 8.50 %
Working Notes:
Cost of new stock is the cost of equity new share which will be issued to raise fresh equity in the company , so it have to incurred flotation cost for issuance of new stock.
Using Gordon growth model : P0 x (1-flotation cost) = D1 / (Ke - g),
ke = Cost of new common stock =??
Flotation cost F =5.00%
Po=current share price = $105.25 per share
g= growth rate= 4.0%
D1= Expected Dividend at t=1=4.50
P0 x (1-flotation cost) = D1 / (Ke - g),
Ke = (D1/(P0 x (1-flotation cost))) + g
Ke = (4.50/(105.25 x (1-0.05))) + 4.0%
Ke = 0.045005626 + 4.0%
Ke = 0.085005626
Ke=8.50056%
Ke = 8.50%
Cost of new stock is 8.50%
Notes: cost of new stock is higher than the cost of retained earnings as for cost of new stock have to incurred flotation cost also.
Please feel free to ask if anything about above solution in comment section of the question.